# Active Adaptive Experimental Design for Treatment Effect Estimation with Covariate Choices

### v1

@@ B.5 Derivation of the Semiparametric Efficiency Bound
$\Bigg{\{}\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\Bigg{\}}\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ +2\Bigg{\{}\theta_{0}(x)-\widehat{\theta}_{t}(x)\Bigg{\}}\frac{q(x)}{\widehat{p}_{t}(x)}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}+\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}$
$\xrightarrow{\mathrm{a.s}}\Bigg{\{}\frac{\big{(}Y_{t}(1)-\mu(1)(x)\big{)}^{2}}{w^{*}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\mu(0)(x)\big{)}^{2}}{w^{*}(0\mid x)}\Bigg{\}}\frac{q^{2}(x)}{p^{*2}(x)}.$
@@
$W_{t}(x)$$\coloneqq\Bigg{\{}\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\Bigg{\}}\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ +2\Bigg{\{}\theta_{0}(x)-\widehat{\theta}_{t}(x)\Bigg{\}}\frac{q(x)}{\widehat{p}_{t}(x)}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}+\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}$
$\ \ \ -\Bigg{\{}\frac{\big{(}Y_{t}(1)-\mu(1)(x)\big{)}^{2}}{w^{*}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\mu(0)(x)\big{)}^{2}}{w^{*}(0\mid x)}\Bigg{\}}\frac{q^{2}(x)}{p^{*2}(x)}-\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}.$
@@
Next, we show $\mathbb{E}_{P_{t}}[W_{t}(X_{t})]\to 0$ by using the $L^{r}$-convergence theorem (Proposition D.5). Note that
$\mathbb{E}_{P_{t}}\left[\Bigg{\{}\frac{\big{(}Y_{t}(1)-\mu(1)(X_{t})\big{)}^{2}}{w^{*}(1\mid X_{t})}-\frac{\big{(}Y_{t}(0)-\mu(0)(X_{t})\big{)}^{2}}{w^{*}(0\mid X_{t})}\Bigg{\}}\frac{q^{2}(X_{t})}{p^{*2}(X_{t})}\right]=\tau^{*}.$
@@
Step 2: check of condition (b).
@@
Step 3: check of condition (c).
@@
=($=\Bigg{(}\Bigg{\{}\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\Bigg{\}}\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$*\displaystyle\ \ \ \ \ \ \ \ \ +2\Bigg{\{}\theta_{0}(x)-\widehat{\theta}_{t}(x)\Bigg{\}}\frac{q(x)}{\widehat{p}_{t}(x)}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}+\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}\Bigg{)}/\tau^{*}$
$\ \ \ -\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\Bigg{\{}\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\Bigg{\}}\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ \ \ \ \ \ \ +2\Bigg{\{}\theta_{0}(x)-\widehat{\theta}_{t}(x)\Bigg{\}}\frac{q(x)}{\widehat{p}_{t}(x)}\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}+\Bigg{\{}\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\Bigg{\}}^{2}\Bigg{)}/\tau^{*}\mid\mathcal{F}_{t-1}\Bigg{]}.$
@@
$\frac{1}{T}\sum^{T}_{t=1}\xi^{2}_{t}-1=\frac{1}{T}\sum^{T}_{t=1}\Bigg{\{}\xi^{2}_{t}-\mathbb{E}\big{[}\xi^{2}_{t}\mid\mathcal{F}_{t-1}\big{]}+\mathbb{E}\big{[}\xi^{2}_{t}\mid\mathcal{F}_{t-1}\big{]}-1\Bigg{\}}\xrightarrow{\mathrm{p}}0.$
Conclusion.
@@
Appendix E Details of the Simulation Studies
@@
E.1 Details of Alternative Experiments
@@
$\widehat{\theta}_{T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=1]\Big{\{}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\Big{\}}}{w^{*}(1\mid X_{t})}-\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=0]\Big{\{}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\Big{\}}}{w^{*}(0\mid X_{t})}+\frac{1}{T}\sum^{T}_{t=1}\mathbb{E}_{X\sim q(x)}\left[\widehat{\mu}_{t}(1)(X)-\widehat{\mu}_{t}(0)(X)\right],$
@@
$\widehat{\theta}_{T}$ $\coloneqq\frac{1}{T}\sum^{T}_{t=1}\left\{\frac{\mathbbm{1}[A_{t}=1]\Big{\{}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\Big{\}}}{w^{*}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\Big{\{}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\Big{\}}}{w^{*}(0\mid X_{t})}\right\}\frac{q(X_{t})}{p^{*}(X_{t})}$
@@
$\widehat{\theta}_{T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=1]\Big{\{}Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\Big{\}}}{w_{t}(1\mid X_{t})}-\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=0]\Big{\{}Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\Big{\}}}{w_{t}(0\mid X_{t})}+\frac{1}{T}\sum^{T}_{t=1}\mathbb{E}_{X\sim q(x)}\left[\widehat{\mu}_{t}(1)(X)-\widehat{\mu}_{t}(0)(X)\right].$
E.2 MSE with Various $T$
@@ E.4 Coverage Ratio
E.5 Simulation Studies with Different Settings
@@
Simulation studies with homogeneous variances.
@@
Simulation studies with homogeneous means.
@@
Appendix F Extensions
@@
F.1 Rejection Sampling from $q(x)$
@@
$\widehat{\theta}_{T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\Psi_{t}(Y_{t},A_{t},X_{t};w_{t},p_{t}),$ (8)
@@
$\widetilde{\Psi}_{t}(Y_{t},A_{t},X_{t};\widehat{w}_{t},\widehat{p}_{t})\coloneqq\Bigg{\{}\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}+\widehat{\theta}_{t}(X_{t})\Bigg{\}}\frac{1}{e_{t}(X_{t})}.$
@@ F.3 Multiple Treatments
For $\theta_{0,K}$, the semiparametric efficiency bound is given as $\tau_{K}(w,p)=\mathbb{E}_{X\sim p(x)}\left[\sum_{a\in[K]}\frac{\sigma^{2}_{0}(a)(X)}{w(a\mid X)}\frac{q^{2}(X)}{p^{2}(X)}\right]$. Because the proof is almost the same as that for Theorem 3.1, we omit the proof.

### v2

@@ A.5 Derivation of the Semiparametric Efficiency Bound
$\left(\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\right)\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ +2\left(\theta_{0}(x)-\widehat{\theta}_{t}(x)\right)\frac{q(x)}{\widehat{p}_{t}(x)}\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)+\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}$
$\xrightarrow{\mathrm{a.s}}\left(\frac{\big{(}Y_{t}(1)-\mu(1)(x)\big{)}^{2}}{w^{*}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\mu(0)(x)\big{)}^{2}}{w^{*}(0\mid x)}\right)\frac{q^{2}(x)}{p^{*2}(x)}.$
@@
$W_{t}(x)$ $\coloneqq\left(\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\right)\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ +2\left(\theta_{0}(x)-\widehat{\theta}_{t}(x)\right)\frac{q(x)}{\widehat{p}_{t}(x)}\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)+\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}$
$\ \ \ -\left(\frac{\big{(}Y_{t}(1)-\mu(1)(x)\big{)}^{2}}{w^{*}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\mu(0)(x)\big{)}^{2}}{w^{*}(0\mid x)}\right)\frac{q^{2}(x)}{p^{*2}(x)}-\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}.$
@@
Next, we show $\mathbb{E}_{P_{t}}[W_{t}(X_{t})]\to 0$ by using the $L^{r}$-convergence theorem (Proposition C.5). Note that
$\mathbb{E}_{P_{t}}\left[\left(\frac{\big{(}Y_{t}(1)-\mu(1)(X_{t})\big{)}^{2}}{w^{*}(1\mid X_{t})}-\frac{\big{(}Y_{t}(0)-\mu(0)(X_{t})\big{)}^{2}}{w^{*}(0\mid X_{t})}\right)\frac{q^{2}(X_{t})}{p^{*2}(X_{t})}\right]=\tau^{*}.$
@@
Step 2: check of condition (b)
@@
Step 3: check of condition (c)
@@
$=\Bigg{(}\left(\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\right)\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ \ \ \ \ \ \ +2\left(\theta_{0}(x)-\widehat{\theta}_{t}(x)\right)\frac{q(x)}{\widehat{p}_{t}(x)}\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)+\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}\Bigg{)}/\tau^{*}$
$\ \ \ -\mathbb{E}_{P_{t}}\Bigg{[}\Bigg{(}\left(\frac{\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(x)\big{)}^{2}}{\widehat{w}_{t}(1\mid x)}-\frac{\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(x)\big{)}^{2}}{\widehat{w}_{t}(0\mid x)}\right)\frac{q^{2}(x)}{\widehat{p}^{2}_{t}(x)}$
$\ \ \ \ \ \ \ \ \ +2\left(\theta_{0}(x)-\widehat{\theta}_{t}(x)\right)\frac{q(x)}{\widehat{p}_{t}(x)}\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)+\left(\mathbb{E}_{\widetilde{X}\sim q(x)}\left[\widehat{\theta}_{t}(\widetilde{X})\right]-\theta_{0}\right)^{2}\Bigg{)}/\tau^{*}\mid\mathcal{F}_{t-1}\Bigg{]}.$
@@
$\frac{1}{T}\sum^{T}_{t=1}\xi^{2}_{t}-1=\frac{1}{T}\sum^{T}_{t=1}\left(\xi^{2}_{t}-\mathbb{E}\big{[}\xi^{2}_{t}\mid\mathcal{F}_{t-1}\big{]}+\mathbb{E}\big{[}\xi^{2}_{t}\mid\mathcal{F}_{t-1}\big{]}-1\right)\xrightarrow{\mathrm{p}}0.$
Conclusion
@@
Appendix D Details of the Simulation Studies
@@
D.1 Details of Alternative Experiments
@@
$\widehat{\theta}_{T}$ $\coloneqq\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=1]\left(Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\right)}{w^{*}(1\mid X_{t})}-\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=0]\left(Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\right)}{w^{*}(0\mid X_{t})}$
$\ \ \ \ \ \ \ \ \ +\frac{1}{T}\sum^{T}_{t=1}\mathbb{E}_{X\sim q(x)}\left[\widehat{\mu}_{t}(1)(X)-\widehat{\mu}_{t}(0)(X)\right],$
@@
$\widehat{\theta}_{T}$ $\coloneqq\frac{1}{T}\sum^{T}_{t=1}\left(\frac{\mathbbm{1}[A_{t}=1]\left(Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\right)}{w^{*}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\left(Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\right)}{w^{*}(0\mid X_{t})}\right)\frac{q(X_{t})}{p^{*}(X_{t})}$
@@
$\widehat{\theta}_{T}$ $\coloneqq\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=1]\left(Y_{t}-\widehat{\mu}_{t}(1)(X_{t})\right)}{w_{t}(1\mid X_{t})}-\frac{1}{T}\sum^{T}_{t=1}\frac{\mathbbm{1}[A_{t}=0]\left(Y_{t}-\widehat{\mu}_{t}(0)(X_{t})\right)}{w_{t}(0\mid X_{t})}$
D.2 MSE with Various $T$
@@ D.4 Coverage Ratio
D.5 Simulation Studies with Different Settings
@@
Simulation studies with homogeneous variances
@@
Simulation studies with homogeneous means
@@
Simulation studies with higher dimensional covariates
[…]
As in the experiment in the main section, we specify the means as $\mathbb{E}_{X\sim q(x)}[Y(1)]=10$ and $\mathbb{E}_{X\sim q(x)}[Y(0)]=7$, under which the ATE is $\theta_{0}=3$. For any $X_{t}$, the conditional means are
$\mu_{0}(1)(X_{t})$ $\coloneqq C_{1}\left(-X_{1,t}+3X_{1,t}^{2}-1\right),$
$\mu_{0}(0)(X_{t})$ $\coloneqq C_{0}\left(0.1X_{1,t}+0.2\right),$
We define $\sigma^{2}_{0}(a)(X_{t})$ as follows:
$\sigma^{2}_{0}(1)(X_{t})$ $\coloneqq 2+1.2\sin(2X_{1,t})+\frac{X_{1,t}+2X^{2}_{1,t}}{25},$
$\sigma^{2}_{0}(0)(X_{t})$ $\coloneqq 2+0.8\cos\left(\frac{X_{1,t}}{2}\right)+\frac{X_{1,t}^{2}}{50}.$
Then, as in Figures 77, we report the empirical MSEs for settings with each dimension $d=3,10,20$. The results are shown in Figure 16.
Figure 16: Results of simulation studies with covariates following $3$-dimensional and $20$-dimensional Gaussian distributions.
D.6 Simulation Studies with the Semi-Synthetic Dataset
This section provides simulation studies using the semi-synthetic dataset called the Infant Health and Development Program (IHDP).
The IHDP dataset has simulated outcomes and real-world covariates (Hill, 2011). The total sample size is $747$, and the covariates consist of $6$ continuous variables and $19$ binary variables.
The outcomes are generated artificially, while the covariates are real-world data. In (Hill, 2011), there are two scenarios called response surface A and response surface B. In response surface A, the potential outcomes $Y_{t}(1)$ and $Y_{t}(0)$ are generated as
$Y_{t}(0)$ $\sim\mathcal{N}(X^{\top}_{t}\bm{\gamma}_{A},1),$
$Y_{t}(1)$ $\sim\mathcal{N}(X^{\top}_{t}\bm{\gamma}_{A}+4,1),$
where each element of $\bm{\gamma}_{A}\in\mathbb{R}^{25}$ is randomly generated from $\{0,1,2,3,4\}$ with probabilities $(0.5,0.2,0.15,0.1,0.05)$.
In response surface B, the potential outcomes $Y_{t}(1)$ and $Y_{t}(0)$ are generated as
$Y_{t}(0)$ $\sim\mathcal{N}\left(\exp\left((X_{t}+W)^{\top}\bm{\gamma}_{B}\right),1\right),$
$Y_{t}(1)$ $\sim\mathcal{N}(X^{\top}_{t}\bm{\gamma}_{B}-q,1),$
where $\bm{W}$ is an offset matrix of the same dimension as $X_{t}$ with every value equal to $0.5$, $q$ is a constant to normalize the average treatment effect conditional on $d=1$ to be $4$, and each element of $\bm{\gamma}_{B}\in\mathbb{R}^{25}$ is randomly generated from $\{0,0.1,0.2,0.3,0.4\}$ with probabilities $(0.6,0.1,0.1,0.1,0.1)$.
Our experiment fixes $T=2000$. In each round $t$, we sample covariate $X_{t}$ with replacement from the $747$ samples. Then, following each setting of the response surfaces, we generate the corresponding outcomes.
As in Figures 77, we report the empirical MSEs for settings with each dimension $d=3,10,20$. The results are shown in Figure 17.
Figure 17: Results of simulation studies with the IHDP dataset (response surface A and B).
Appendix E Extensions
@@
E.1 Rejection Sampling from $q(x)$
@@
$\widehat{\theta}_{T}\coloneqq\frac{1}{T}\sum^{T}_{t=1}\widetilde{\Psi}_{t}(Y_{t},A_{t},X_{t};w_{t},p_{t}),$ (8)
@@
$\widetilde{\Psi}_{t}(Y_{t},A_{t},X_{t};\widehat{w}_{t},\widehat{p}_{t})\coloneqq\left(\frac{\mathbbm{1}[A_{t}=1]\big{(}Y_{t}(1)-\widehat{\mu}_{t}(1)(X_{t})\big{)}}{\widehat{w}_{t}(1\mid X_{t})}-\frac{\mathbbm{1}[A_{t}=0]\big{(}Y_{t}(0)-\widehat{\mu}_{t}(0)(X_{t})\big{)}}{\widehat{w}_{t}(0\mid X_{t})}+\widehat{\theta}_{t}(X_{t})\right)\frac{1}{e_{t}(X_{t})}.$
@@ E.3 Multiple Treatments
For $\theta_{0,K}$, the semiparametric efficiency bound is given as $\tau_{K}(w,p)=\mathbb{E}_{X\sim p(x)}\left[\sum_{a\in[K]}\frac{\sigma^{2}_{0}(a)(X)}{w(a\mid X)}\frac{q^{2}(X)}{p^{2}(X)}\right]=\mathbb{E}_{X\sim q(x)}\left[\sum_{a\in[K]}\frac{\sigma^{2}_{0}(a)(X)}{w(a\mid X)}\frac{q(X)}{p(X)}\right]$. Because the proof is almost the same as that for Theorem 3.1, we omit the proof.